Ask your own question, for FREE!
Mathematics 8 Online
OpenStudy (anonymous):

A bag contains 2 white marbles and 7 purple marbles. Two marbles are drawn at random. One marble is drawn and not replaced. Then a second marble is drawn. What is the probability of selecting a purple marble and then a white marble? What is the probability of selecting two white marbles? Is there a greater chance of selecting two white marbles in a row or two purple marbles in a row?

OpenStudy (anonymous):

Please help me! Thanks. -Vivid

OpenStudy (madig):

1) \[7/9*2/8= 7/36\] 2) \[2/9*1/8=1/36\] 3) 2 white marbles: \[2/9*1/8=1/36\] 2 purple marbles: \[7/9*6/8= 7/12\]

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!